OpenStudy (anonymous):

I'm trying to solve this matrix but need help. 1 2 -9 -8 0 1 7 -3 0 0 1 7 my choice of answers are; a -2, -11, 6 b 159, -52, 7 c -8, -3, 7 d 21, 46, 7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

im guessing b at sight

OpenStudy (anonymous):

what do you mean "solve" it? do you mean find the determinant? do you mean find the eigenvalues?

OpenStudy (anonymous):

oh, i see. a system of equations.

OpenStudy (amistre64):

you pretty much work it up from the bottom; that 1,7 means teh last value is = to 7

OpenStudy (amistre64):

1 7 , -3 means 1 +7(7) = -3; the second so 1 = -52

OpenStudy (amistre64):

1 2 -8, -8 then means: 1 +2(-52) -9(7) = -8

OpenStudy (amistre64):

and i bet that ends up 1 = 159

OpenStudy (anonymous):

Thanks, this stuff is hard for this old mind to get a handle on.

OpenStudy (amistre64):

\[\begin{array}c x&y&z&|&=\\ 1&2& -9&|& -8\\ 0& 1& 7& |&-3\\ 0& 0& 1& |&7 \end{array}\]